OpenStudy (anonymous):

17% of what number is 0.5? round to nearest tenth.

OpenStudy (campbell_st):

if 17% = 0.5 find 1% by dividing both sides by 17 then multiply both sides by 100 to get 100% the number you need.

OpenStudy (anonymous):

17% divided by 0.5 i got 0.02941?

OpenStudy (anonymous):

.17x=0.5 x=0.5/.17

OpenStudy (anonymous):

0.5/.17=50/17 then a calculator

OpenStudy (anonymous):

im still confused, i still need to show the work but what exactly am i looking for in this problem?

OpenStudy (anonymous):

0.085?

OpenStudy (campbell_st):

if 0.5/17 = 0.02941 then multiply by 100 next 0.02941 x 100 =

OpenStudy (anonymous):

i gotta try it again, 0.02941 was wrong, but to round it to the nearest tenth, since its under 5 i could put it as .0294 but its saying my full math is wrong

OpenStudy (anonymous):

what answer am i looking for? i still have to do the math but what am i looking for?

OpenStudy (campbell_st):

do the rounding when you have your answer... not before

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1 sec

OpenStudy (campbell_st):

ok 0.5/17 = 0.02941 so 1% = 0.02941 multiply by 100 to get 100% of the number....

OpenStudy (campbell_st):

so your answer should be 2.941 rounding to the nearest 10th or 1 decimal place is 2.9 to check calculate 2.9 x 0.17 =

OpenStudy (anonymous):

got it thanks